OpenStudy (anonymous):

K im confused..... |2x + 5| = x + 1 x= -4 OR x= -2

OpenStudy (xapproachesinfinity):

Wrong

OpenStudy (anonymous):

If I'm doing |2(-4) + 5| doesnt this equal |-3| and it becomes a positive 3?

OpenStudy (xapproachesinfinity):

many people today are asking absolute values question what's going on. my head is going dizzy with this hehehe

OpenStudy (xapproachesinfinity):

i don't what you are trying to do here where did you come up with x=-4 or x=-2 did you solve the equation first?

OpenStudy (anonymous):

You solve for it first... and then you plug it in

OpenStudy (swissgirl):

ummm to me x=-4 and x-2 both wont work but i may be wrong

OpenStudy (xapproachesinfinity):

okay then you must have done it wrong, because the two answer are not good

OpenStudy (anonymous):

no thats the right answer when you solve for x

OpenStudy (anonymous):

i already know it doesnt work but i need to know the process

OpenStudy (xapproachesinfinity):

no it is not the right answer! check you steps

OpenStudy (xapproachesinfinity):

hey don't disrespect people, respect yourself pls

OpenStudy (anonymous):

i'm confused

OpenStudy (xapproachesinfinity):

so @yomamabf show us the work you did

OpenStudy (anonymous):

@swissgirl what do you mean by x + 1 is greater than or equal to 0

OpenStudy (anonymous):

wait can you just tell me if its possible though? if |2x + 5| equals positive 3 if x is -4

OpenStudy (swissgirl):

yes

OpenStudy (swissgirl):

and that was an observation i made basically x must equal or be greater than -1

OpenStudy (anonymous):

why -1?

OpenStudy (anonymous):

oh okay so i was right it is positive 3 because of the absolute value

OpenStudy (swissgirl):

ehhh i know how you got -4 and -2

OpenStudy (anonymous):

yea i solved for x

OpenStudy (swissgirl):

yup but for some reason when you plug it back in it ... it doesnt add up

OpenStudy (anonymous):

yea its not supposed to i just need to understand why

OpenStudy (swissgirl):

most likely cuz of the x on the other side

OpenStudy (anonymous):

hey it doesnt equal each other because its a 3 = -3? right?

OpenStudy (swissgirl):

Yup :)

OpenStudy (anonymous):

thanks love

OpenStudy (swissgirl):

no problem hun

OpenStudy (anonymous):

if i plug in -2 to -x-1 how does it come out to be -1 i'm confused

OpenStudy (anonymous):

oh oops nvm

ganeshie8 (ganeshie8):

are you solving |2x + 5| = x + 1 ?

OpenStudy (anonymous):

no nvm i got it thanx

ganeshie8 (ganeshie8):

No worries! I am not sure how you have worked it but squaring both sides and solving the quadratic seems to be one way to approach this. Also don't forget to check your solutions everytime you square because, squaring increases the degree of polynomial and the end result is increase in number of fake solutions (I see you're checking your answers already :))

OpenStudy (swissgirl):

we didnt solve it ... we just showed how x=-2 and x=-4 will not satisfy the equations. Are there any other options

OpenStudy (swissgirl):

Like any other methods of solving this absolute value equation?

ganeshie8 (ganeshie8):

I like squaring both sides because I don't have to worry about the messy cases and + and - signs

OpenStudy (swissgirl):

Thanks :)

ganeshie8 (ganeshie8):

Without squaring : |2x + 5| = x + 1 gives two cases : 2x+5 = x+1 OR 2x+5 = -(x+1) x = -4 OR x = -2

ganeshie8 (ganeshie8):

not bad at all !

OpenStudy (anonymous):

thanks a lot for your help!

ganeshie8 (ganeshie8):

however both are extraneous solutions because when you plug them back in the original equaiton, they don't satisfy

ganeshie8 (ganeshie8):

don't mean to be repetitive, just saying for completeness :)

OpenStudy (anonymous):

yap gotcha thanks!

OpenStudy (swissgirl):

hahaha :D